This isn't a test or textbook question. Here is the question: Mary bowled 199 and this raised her average from 177 to 178. To raise her average to 179 with the next game, what must she bowl?
Quick help would be appreciated, thanks a lot,
Cameron.
: This isn't a test or textbook question. Here is the question: Mary bowled 199 and this raised her average from 177 to 178. To raise her average to 179 with the next game, what must she bowl?
Quick help would be appreciated, thanks a lot,
Cameron.

Answer by scott8148(2801)

the average is the sum of her scores (S) divided by the number of scores (n)

initially, S/n=177 __ S=177n

after 199, (S+199)/(n+1)=178 __ S+199=178n+178 __ substituting __ 177n+199=178n+178

subtracting 177n+178 __ 21=n

in order to raise her average one point, she scored 22 points above average in her 22nd game

to raise her average another point, she will have to score 23 points above her average in her 23rd game

178+23=201
Question 155211: This isn't a test or textbook question. Here is the question: Mary bowled 199 and this raised her average from 177 to 178. To raise her average to 179 with the next game, what must she bowl?
Quick help would be appreciated, thanks a lot,
Cameron.
: This isn't a test or textbook question. Here is the question: Mary bowled 199 and this raised her average from 177 to 178. To raise her average to 179 with the next game, what must she bowl?
Quick help would be appreciated, thanks a lot,
Cameron.

Answer by edjones(2415)

Let x=number of games she bowled.
Average=total scores of all games bowled / x
178(x+1)=177x+199
178x+178=177x+199
x=21 number of games bowled to get 177 average.
The 199 score was game #22
.
Let y=the game score required to raise her average to 179.
(178*22)+y=179*23
3916+y=4117
y=201
.
Ed